# Green-NAS: Efficient Neural Architecture Search for On-Device Weather Forecasting

### v1

@@ Abstract
We introduce Green-NAS, a multi-objective NAS (neural architecture search) framework designed for low-resource environments using weather forecasting as a case study. By adhering to ‘Green AI’ principles, the framework explicitly minimizes computational energy costs and carbon footprints, prioritizing sustainable deployment over raw computational scale. The Green-NAS architecture search method is optimized for both model accuracy and efficiency to find lightweight models with high accuracy and very few model parameters; this is accomplished through an optimization process that simultaneously optimizes multiple objectives. Our best-performing model, Green-NAS-A, achieved an RMSE of 0.0988 (i.e., within 1.4% of our manually tuned baseline) using only 153k model parameters, which is 239 times fewer than other globally applied weather forecasting models, such as GraphCast. In addition, we also describe how the use of transfer learning will improve the weather forecasting accuracy by approximately 5.2%, in comparison to a naive approach of training a new model for each city, when there is limited historical weather data available for that city.
@@ VII Conclusion
Code and trained models will be made publicly available upon acceptance. Historical weather data was obtained from the publicly accessible Open-Meteo Historical Weather API (https://open-meteo.com).

### v2

# Green-NAS: A Global-Scale Multi-Objective Neural Architecture Search for Robust and Efficient Edge-Native Weather Forecasting

@@ Abstract
We introduce Green-NAS, a multi-objective neural architecture search (NAS) framework designed for low-resource environments using weather forecasting as a case study. By adhering to ‘Green AI’ principles, the framework explicitly minimizes computational energy costs and carbon footprints, prioritizing sustainable deployment over raw computational scale. The Green-NAS architecture search method is optimized for both model accuracy and efficiency to find lightweight models with high accuracy and very few model parameters; this is accomplished through an optimization process that simultaneously optimizes multiple objectives. Our best-performing model, Green-NAS-A, achieved an RMSE of 0.0988 (i.e., within 1.4% of our manually tuned baseline) using only 153K model parameters, which is 239 times fewer than other globally applied weather forecasting models, such as GraphCast. In addition, we also describe how the use of transfer learning improves the weather forecasting accuracy by approximately 5.2%, in comparison to a naive approach of training a new model for each city, when there is limited historical weather data available for that city.
@@ V-B Performance Comparison
Table I presents a comprehensive comparison of all evaluated models, including our NAS-discovered architectures, manually tuned baselines, and the state-of-the-art GraphCast model. Our Green-NAS models achieve competitive accuracy while being orders of magnitude smaller, demonstrating the effectiveness of multi-objective NAS for edge deployment.
TABLE I: Comparison of Green-NAS Models with Baselines and State-of-the-Art
Model Params RMSE Inf. (ms) Size
Statistical Baselines
Persistence 0 0.1842
Climatology 0 0.1523
Manually Designed DL Models
Manual GRU 153K 0.1004 0.29 598 KB
Manual Transformer 135K 0.0974 1.89 527 KB
State-of-the-Art (Global Scale)
GraphCast [30] 36.7M $\sim$1 GB
Green-NAS (Ours)
Green-NAS-A 153K 0.0988 0.47 598 KB
Green-NAS-B 4.2K 0.0996 0.29 17 KB
Green-NAS-C 1.1K 0.1019 0.33 4 KB
Not directly comparable; GraphCast solves global gridded forecasting.
@@
Figure 2: Architecture Comparison. Performance comparison across three dimensions: (A) Forecast accuracy (RMSE), (B) Model size (parameters, log scale), and (C) Inference latency. Green-NAS models show competitive accuracy with superior efficiency compared to manual baselines.
@@ V-C Transfer Learning Efficacy
Table II presents a detailed breakdown of transfer learning performance across varying data availability levels. At every data fraction, transfer learning consistently outperforms training from scratch with high statistical significance.
TABLE II: Transfer Learning vs. Training From Scratch at Varying Data Fractions
Data Samples Scratch Transfer Impr. p-value
1% 3,155 0.1083 0.1020 +5.7% $10^{-8}$
10% 31,550 0.1016 0.0980 +3.5% $10^{-8}$
50% 157,752 0.0967 0.0921 +4.7% $10^{-14}$
100% 315,504 0.0948 0.0898 +5.2% $10^{-12}$
@@ VII Conclusion
Code and trained models are publicly available at https://github.com/Muhtasim-Munif-Fahim/Green-NAS. Historical weather data was obtained from the publicly accessible Open-Meteo Historical Weather API (https://open-meteo.com).